OpenStudy (anonymous):

PPPPLLLLLZZZZZ HHHHEEELLLPPPP!!!!!!!! giving out a medal what is the product of 2√u(√5+3√45)

OpenStudy (anonymous):

do you mean 20√5u

OpenStudy (anonymous):

because ur answer is not evan an option

OpenStudy (anonymous):

Disregard my previous solution.

OpenStudy (anonymous):

ok then what is it?

OpenStudy (anonymous):

I assume you want to find the result of the following: \[2 \sqrt{u} (\sqrt{5}+3\sqrt{45})\] right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

id also like if u could explain to me how u got it if thats alright with you

OpenStudy (anonymous):

\[\sqrt{45} = \sqrt{9*5} = \sqrt{9}*\sqrt{5} = 3\sqrt{5}\]

OpenStudy (anonymous):

thats not one of the anwsers options though :(

OpenStudy (anonymous):

Wait please, I am trying to explain each one of the terms. Do you understand till now?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

You have the square root of 45 in your expression. We have simplified that above to 3 times square root of 5. Do you understand that?

OpenStudy (anonymous):

Sort of, u do the factor tree of 45 and your prime numbers will be 5,3,and 3 since there are two 3's one goes on the outside right?

OpenStudy (anonymous):

Yeah, we have 3^2 in the square root, so it can be a 3 outside it

OpenStudy (anonymous):

ok i got that part but what about the rest of the equation

OpenStudy (anonymous):

So here it is. We said that \[\sqrt{45} = 3\sqrt{5}\] So we would have: \[2\sqrt{u}(\sqrt{5}+3\sqrt{45}) = 2\sqrt{u}(\sqrt{5}+3*3\sqrt{5}) = 2\sqrt{u}(\sqrt{5}+9\sqrt{5})\]

OpenStudy (anonymous):

When you read this, tell me if you understand, so I will continue.

OpenStudy (anonymous):

ok i think i understand all u do is multiply the outside 3 by the 3 being added to the 5 which gives you 9

OpenStudy (anonymous):

All we did, is to replace the square root of 45, with 3 square root of 5, that we have found before and do the 3 times 3 multiplication

OpenStudy (anonymous):

ok i got it

OpenStudy (anonymous):

May I ask you know, if we add: \[\sqrt{5} + 9\sqrt{5}\] what would be the result?

OpenStudy (anonymous):

would it be 9√10????

OpenStudy (anonymous):

No. We have: \[\sqrt{5}+9\sqrt{5} = 1\sqrt{5}+9\sqrt{5}\] Let's say \[\sqrt{5}\] is an apple. So we would have: \[1 apple + 9 apples = ?\]

OpenStudy (anonymous):

10 apples xD

OpenStudy (anonymous):

Right. So 1 square root of 5 plus 9 square roots of 5 would be ?

OpenStudy (anonymous):

10√10

OpenStudy (anonymous):

or 10√5 because 5 are like term... one of those

OpenStudy (anonymous):

Right! We said square root of 5 is an apple and we would have 10 apples. So we would have 10 square roots of 5! That is: \[\sqrt{5}+9\sqrt{5} = 10 \sqrt{5}\]

OpenStudy (anonymous):

ok got it

OpenStudy (anonymous):

So we can continue with the expression: \[2\sqrt{u}(\sqrt{5}+9\sqrt{5}) = 2\sqrt{u}(10\sqrt{5}) = 2*10*\sqrt{u}\sqrt{5}\] Ok with that?

OpenStudy (anonymous):

so 20√u√5 would be the awnser right?

OpenStudy (anonymous):

Right!

OpenStudy (anonymous):

TY so much could u help me with another one though, if u want i can put another question up and u reply to that one so u can get another metal

OpenStudy (anonymous):

I am sorry, but I will have to log off now. I am sure the other members will help you though! Have fun!

OpenStudy (anonymous):

medal*

OpenStudy (anonymous):

thnx anyway